800 students have a mean weight of 75 kg and standard deviation of 5kg. How many students weigh more than 80 kg


1
Expert's answer
2021-02-03T15:25:28-0500

Let "X" be a random variable such that weight of the students, which is normally distributed with mean "75" kg and standard deviation "5" kg.

Then we have "\mu=75" and "\sigma =5"

Let us consider "Z=\frac{X-\mu}{\sigma }". Then "Z=\frac{X-75}{5}"

Now we have to find "P(X>80)"

"\therefore P(X>80)=P(Z>\frac{80-75}{5})=P(Z>1)=0.5-P(0<Z<1)"

"=0.5-0.3413" "[P(0<Z<1)=0.3413" from normal distribution table "]" "=0.1587"


"\therefore" Number of students having weight more than "80" kg is "=80\u00d7P(X>80)"

"=80\u00d70.1587"

"=127" (approximately)

Which is the required answer.